OpenStudy (anonymous):

Somebody please help me!!! Will fan and medal! I have been here for 20 minutes and never received any help! The table below represents a linear function f(x) and the equation represents a function g(x): x f(x) g(x) -1 -5 g(x) = 4x + 3 0 -1 1 3 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). Part B: Which function has a greater y-intercept? Justify your answer.

OpenStudy (amistre64):

part B is simple, what value of x do we need to find any point on the y axis?

OpenStudy (amistre64):

that sentence is bad lol .... what value of x do we need in order to determine any point on the y axis;

OpenStudy (anonymous):

Oh god, I don't know. I am usually good with functions but today I'm so confused...

OpenStudy (anonymous):

to find the slope you have to do y2-y1/x2-x1

OpenStudy (amistre64):

you do agree that the x axis is numbered, positive xs values to the right, and negative to the left of the y axis, correct?

OpenStudy (amistre64):

the y axis actually intercepts the x axis at the value x=??

OpenStudy (anonymous):

I only got the explaination for g(x) = 4x + 3 it's x f(x) -1 -5 0 -1 1 3 That is what I need help with...

OpenStudy (amistre64):

find g(-1), g(0), and g(1) to fill in your table

OpenStudy (anonymous):

g(-1) = 4x + 3 4 * -1 = -4 -4 + 3 = -1

OpenStudy (anonymous):

0=x1 -1=x2 -1=y1 -5=y2

OpenStudy (amistre64):

good, g(-1) = 4(-1)+3 = -1 is correct for that one

OpenStudy (anonymous):

so -5-(-1)/-1-0 will be your slope for f(x)

OpenStudy (anonymous):

g(0) = 4x + 3 4 * 0 = 0 0 + 3 = 3

OpenStudy (amistre64):

youre doing fine

OpenStudy (anonymous):

y=mx+b m=slope g(x)=4x+3 what is the slope for g(x)

OpenStudy (anonymous):

g(1) = 4x + 3 4 * 1 = 4 4 + 3 = 7

OpenStudy (amistre64):

very good

OpenStudy (anonymous):

I already got the explanation for g(x) Functions like f(x), g(x), etc. are equal to y. g(x) = 4x + 3 y = 4x + 3 y = mx + b m = slope b = y-intercept m = 4 b = 3 The slope of the equation g(x) = 4x + 3 is 4.

OpenStudy (anonymous):

ok, so do you know the slope for f(x)

OpenStudy (amistre64):

since the y axis crosses the x axis at x=0; the y intercept of any function will be deterined when x=0

OpenStudy (anonymous):

nope, that is what i am trying to figure out...

OpenStudy (anonymous):

to find the slope you have to do y2-y1/x2-x1 0=x1 -1=x2 -1=y1 -5=y2 so -5-(-1)/-1-0 will be your slope for f(x)

OpenStudy (amistre64):

its not stated to clearly, but a function like f(x) is an object that can be measured by comparing its distance and direction from the x axis at a given value of x. the function, if it has a y intercept, will cross over the y axis when x=0

OpenStudy (anonymous):

wait don't you mean y^2 - y^1/x^2 - x^1

OpenStudy (amistre64):

no, subscripts are not exponents

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh ok, lol sorry,

OpenStudy (anonymous):

\[\frac{ y _{2} -y _{1}}{ x _{2} -x _{1}}\]

OpenStudy (anonymous):

I know that this may not be correct but I got 4.5 .....

OpenStudy (anonymous):

-5-(-1)=-5+1=??

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

-1-0=??

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

-4/-1=??

OpenStudy (anonymous):

4

OpenStudy (anonymous):

so your slope for f(x) is 4

OpenStudy (anonymous):

Okay, thanks.. Sorry it's just the the lesson did a terrible job of explaining...

OpenStudy (anonymous):

wait, what about (1,3)

OpenStudy (anonymous):

its ok :)

OpenStudy (anonymous):

you don't need to use it

OpenStudy (anonymous):

you only need two coordinates to find slope

OpenStudy (anonymous):

Ok,

OpenStudy (anonymous):

so now you need to find the y intercept of f(x)

OpenStudy (anonymous):

y=mx+b b=y intercept

OpenStudy (anonymous):

we know that so far your equation looks like: y=4x+b

OpenStudy (anonymous):

y = 4x + b (let's use -5 maybe?) y = 4x + -5 4 * -1 = -4 -4 + (-5) = -9

OpenStudy (anonymous):

no we don't plug anything in for b

OpenStudy (anonymous):

lets use coordinate (1,3)

OpenStudy (anonymous):

1=x and 3=y so the equation would look like: 3=4(1)+b

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

get b by itself and that will by your y intercept for f(x)

OpenStudy (anonymous):

y = mx + b y = 4x + b 4 * 1 = 4 4 + 3 = 7

OpenStudy (anonymous):